OpenStudy (anonymous):

balanced the equation C7H8 + O2 → balanced the equation AgClO3 + SrI2 → balanced the equation Pb(NO3)2 + K2CO3 →

OpenStudy (anonymous):

@Gabebro13

OpenStudy (anonymous):

O_O Sorry, I don't know this... Chemistry isn't really my best subject. But if you need help with biology, just tag me like you did here.

OpenStudy (anonymous):

could you get someone to help me with this

OpenStudy (anonymous):

Sure. @Abhisar

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (anonymous):

thanks

OpenStudy (abhisar):

so u r not given the product side of the reaction ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Predict the products for each of the reactions below. Write a complete balanced equation and enter it in the answer blank.

OpenStudy (abhisar):

ok ! Have u read the reactio of hydrocarbon with oxygen ?

OpenStudy (abhisar):

*reaction

OpenStudy (anonymous):

no

OpenStudy (abhisar):

Hydrocarbon + Oxygen ----> Carbon dioxide + Water. Can u write down the products for first reaction ?

OpenStudy (anonymous):

how

OpenStudy (anonymous):

is it like this h+o->cd+w

OpenStudy (abhisar):

\(\sf C_7H_8+O_2--->CO_2+H_2O\)

OpenStudy (abhisar):

Can u balance it now ?

OpenStudy (abhisar):

????

OpenStudy (anonymous):

yes

OpenStudy (abhisar):

I am always here dear....i was waiting for ur answer...meanwhile helping others ! :)

OpenStudy (abhisar):

so balance it !

OpenStudy (anonymous):

PB(NO3)2(s) + K2CO3(s) = PBCO3(s) + KNO3(s)

OpenStudy (abhisar):

Pb(No3)2(s) + K2Co3(s) = PbCo3(s) + 2 KNo3(s)

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

which answer is that ?is it the first question

OpenStudy (abhisar):

C7H8O2 + 8 O2 = 7 CO2 + 4 H2O

OpenStudy (abhisar):

^^^^ This is for first one.

OpenStudy (abhisar):

Pb(No3)2(s) + K2Co3(s) = PbCo3(s) + 2 KNo3(s) last one

OpenStudy (abhisar):

I am leaving second one upon u.

OpenStudy (anonymous):

the last one is incorrect

OpenStudy (abhisar):

this mean u wrote the equation wrong ! u talking about this one na ?? \(\color{blue}{\text{Originally Posted by}}\) @Abhisar Pb(No3)2(s) + K2Co3(s) = PbCo3(s) + 2 KNo3(s) last one \(\color{blue}{\text{End of Quote}}\)